OpenStudy (anonymous):

y varies inversely with twice the value of x. When y = 10, x = 3. Find y when x = 5.

myininaya (myininaya):

inversely think divide y=k/(2 x)

OpenStudy (anonymous):

Y=K/X inverse, Y = KX direct

OpenStudy (anonymous):

to find the constant of variation (k) we use y = k/x we know that y = 10 when x = 3 10 = k/3 10*3=k = 30 when x is 5, y is: y = k/x y = 30/5 y= 6

OpenStudy (vishal_kothari):

You need to put the first sentence into a line equation, form of y = mx + b. We know the slope is -2, as y goes down by 2 when x goes up by 1 (m = (change in y/change in x). So... y = -2x + b Now, let's solve for b by inputting your known coordinates, (3, 10) 10 = -2 * 3 + b b = 16 So... now we know the full line equation: y = -2x + 16 We can now substitute 5 in for x to find our y value. y = -2(5) + 16 y = -10 + 16 y = 6

myininaya (myininaya):

\[y=\frac{k}{2 x}\] \[10=\frac{k}{2 \cdot 3} => 10=\frac{k}{6} => 60=k\] \[y=\frac{60}{2 x}=\frac{30}{x}\] \[y=\frac{30}{x}\] \[y=\frac{30}{5}=6\]